OpenStudy (anonymous):

Hello! How do I figure out if sigma n=1 to infinity cos(n*pi)/n converges or diverges?

OpenStudy (amistre64):

maybe convert it to its taylor representation?

OpenStudy (amistre64):

\[cos(u)=1+\sum_{k=1}^{\infty}\frac {(-1)^{k}}{(2k)!}u^{2k}\] \[cos(\pi~n)=1+\sum_{1}^{\infty}\frac {(-1)^{n}}{(2k)!}(\pi)^{2k}~(n)^{2k}\] \[\frac{cos(\pi~n)}{n}=1+\sum_{1}^{\infty}\frac {(-1)^{n}}{(2k)!}(\pi)^{2k}~(n)^{2k-1}\] or maybe a ratio test?

OpenStudy (amistre64):

my first ideas bad .... need more MtDew this early

OpenStudy (anonymous):

hmm our teacher never taught us the Taylor representation, I remember her telling us she wasn't going to.. I'll try the ratio test..

OpenStudy (amistre64):

yeah, ratio will help to define the interval of convergence ... at least thats what by addled brain is telling me

OpenStudy (mathmale):

sigma n=1 to infinity cos(n*pi)/n : My first impulse would be to do a comparison test. If you think about it, you'll likely conclude, as I have, that \[\frac{ \cos n \pi }{ n }\le \frac{ 1 }{ n }\]

OpenStudy (mathmale):

Note that the series 1/n is the "harmonic series" and is divergent. Unfortunately, that doesn't help us much in determining whether or not the given series is convergent or divergent.

OpenStudy (amistre64):

alternating series might be useful since we are dealing with an alternating trig

OpenStudy (mathmale):

@amistre64 : an alternating series would include a factor such as (-1)^n, which cos (n*pi) does not have. Aha! inspiration! Would one of you guys please write out the first few terms of the sequence cos n*Pi? let n=0, 1, 2, 3, ... What could you conclude from that?

OpenStudy (amistre64):

cos(pi) = -1 cos(2pi)/2 = 1/2 cos(3pi)/3 = -1/3 cos(4pi)/4 = 1/4

OpenStudy (amistre64):

so it looks to be an alternating hyperG :)

OpenStudy (amistre64):

cos(pi*n) is equal to (-1)^n

OpenStudy (mathmale):

That's a significant discovery. @sgttris @amistre64 Ask yourself: would this alternating series be convergent or divergent, and why?

OpenStudy (amistre64):

im working on the why part at the moment :)

OpenStudy (amistre64):

my own mind goes this far \[\sum_1\frac1{2n}+\sum_1\frac{-1}{2n-1}\] \[\sum_1\frac1{2n}-\frac{1}{2n-1}\] \[\sum_1\frac{2n-2}{2n(2n-1)}\] \[\sum_1\frac{n-1}{n(2n-1)}\]

OpenStudy (amistre64):

ugh, thats off for some reason

OpenStudy (mathmale):

Please look up "alternating series" and refresh your memory regarding the criteria for convergence of an "alt series."

OpenStudy (amistre64):

if lim to 0, and sequence is decreasing, then convergent seems to be the criteria

OpenStudy (amistre64):

since limit of 1/n is 0, and 1/n is decreasing ... still want to see it a different way tho, i like things to be difficult

OpenStudy (amistre64):

i also forgot how to add fraction ... \[\sum_1\frac{2n-1-2n}{2n(2n-1)}\] \[\sum_1\frac{-1}{2n(2n-1)}\] thats better

OpenStudy (mathmale):

We're trying to determine whether the series (-1)^n / n converges or diverges, and recognize this as an alternating series. Separate the (-1)^n from the rest, focusing on \[a _{n}=1/n\] If: (1) this a-sub-n is decreasing with n (2) its limit as n goes to infinity is zero and (3) it's always positive, then we conclude tht the given alternating series converges.

OpenStudy (amistre64):

hmm, the 3rd crit is new to me

OpenStudy (mathmale):

Yes, I was "winging it." How about this: YOU look up "alternating series" and determine more accurately than I what the third criterion is? :)

OpenStudy (amistre64):

yours resembles Leibnez thrm for alternating series. according to math24.net

OpenStudy (amistre64):

i cant seem to find it on Pauls site ... which is usually pretty good about that stuff :/

OpenStudy (amistre64):

Paul seems to be saying that as long as |an| eventually conforms, then its convergent .... shich seems to me to imply that it can be wonky at first but evens out in the end

OpenStudy (anonymous):

idk how to type it and I'm in a rush but if A sub n+1 is less than or equal to A sub n and the lim of n to infinity is =0 then it will converge

OpenStudy (anonymous):

*lim of n to infinity for A sub n

OpenStudy (mathmale):

Yes, that was one of the criteria for convergence: the limit as n goes to inf. of a-sub-n must be zero. I see that both of you know enough about this material so that a quick ref to material on "alternating series" will give you all 3 criteria. You interested enough to check that out (gasp, in a book)? :)

OpenStudy (amistre64):

my archie and jughead doesnt have a chapter on it ;)

OpenStudy (anonymous):

hm, I was only taught that, and I wish I had a book lol, thank you though, the more I know the better

OpenStudy (mathmale):

@amistre64: Don't depend upon Archie or Jughead. Both are numbskulls. Ask Veronica instead.

OpenStudy (mathmale):

@sgttris: I surely appreciate your interest in this material.